OpenStudy (anonymous):

Express the ratio 15 out of 25 as a fraction in simplest form. A. 5/25 B. 15/25 C. 3/5 D. 1/25

OpenStudy (anonymous):

c

OpenStudy (anonymous):

15/25---so what is a common factor of both numer and denom--5 , right? well how many 5's are in each?

OpenStudy (anonymous):

forgot one choice srry

OpenStudy (anonymous):

explain toady

OpenStudy (apoorvk):

\[\frac{15}{25} = \frac{5 \times 3}{5 \times 5 }= \frac{\cancel5 \times 3}{\cancel5 \times 5 }=??\]

OpenStudy (apoorvk):

helps? good that you want to understand, @partyrainbow276

OpenStudy (anonymous):

so..... C

OpenStudy (apoorvk):

yeah! correct-o!

OpenStudy (anonymous):

answer is c

OpenStudy (anonymous):

? wait thats wrong

OpenStudy (anonymous):

its 3/5

OpenStudy (apoorvk):

Yeah, I said you're right - C. is 3/5 right?

OpenStudy (apoorvk):

I meant 'correct-o' - way of saying lol. not "correct-zero"

OpenStudy (anonymous):

o yea my bad!

OpenStudy (apoorvk):

lol it's okay.